OpenStudy (anonymous):

What is the exact value of:

OpenStudy (anonymous):

\[\sqrt{27/64}\]

OpenStudy (anonymous):

.-.

OpenStudy (luigi0210):

Here this will make it easier: \[\sqrt{\frac{ 3*3*3 }{ 8*8 }}\]

OpenStudy (anonymous):

Ohhh snap Luigi, fanned you now your 80 :P

OpenStudy (luigi0210):

Ha, thank you :D

OpenStudy (anonymous):

Welcome c:

OpenStudy (anonymous):

Back to the problem, that does make it a tiny easier; i'm not really good at these.-.

OpenStudy (luigi0210):

Well do you know that sqrt(9)=3 right? and sprt(3*3)=3 is the same

OpenStudy (anonymous):

\[9 \sqrt{3}\] ----------- ? 8

OpenStudy (luigi0210):

So close >.<

OpenStudy (anonymous):

\[3\sqrt{3}\] ----------- ? 8 @Luigi0210

OpenStudy (luigi0210):

There you go bud :)

OpenStudy (anonymous):

You're likable now :P joking you always were xD

OpenStudy (luigi0210):

btw when you take the squareroot of something it's always +/-

OpenStudy (anonymous):

Can you help me with another one?.

OpenStudy (luigi0210):

Sure

OpenStudy (anonymous):

\[\sqrt{50} / 3\sqrt{15}\] exact value

OpenStudy (luigi0210):

Multiply the top and bottom by 3sqrt(15)

OpenStudy (anonymous):

\[25\sqrt{30} / 45 \] ?

OpenStudy (luigi0210):

No, try again

OpenStudy (anonymous):

Hmmmm, maybe \[3\sqrt{30} / 5\] ehh I'm not good at these at all.

OpenStudy (luigi0210):

\[\frac{ \sqrt50 }{ 3\sqrt15 }*\frac{ 3\sqrt15 }{ 3\sqrt15 }\] \[\frac{ 3*\sqrt{50*15} }{ 3*3*\sqrt{15*15} }\]

OpenStudy (anonymous):

I have no idea.

OpenStudy (luigi0210):

At least try brah D:

OpenStudy (anonymous):

\[\sqrt{30} / 9\] \[5\sqrt{30}\] idk something .-.

OpenStudy (luigi0210):

xD

OpenStudy (anonymous):

Is it one of those?.-.

OpenStudy (luigi0210):

Multiply them out

OpenStudy (anonymous):

I dont know how.-. lol.

OpenStudy (luigi0210):

Numbers aren't scary, they're your friend

OpenStudy (anonymous):

I'm so confused. I don't know man.-.

OpenStudy (luigi0210):

Well, where are you confused?

OpenStudy (anonymous):

The multiplying part.

OpenStudy (luigi0210):

Well, according to the math rules: you are not allowed to have a radical in denominator.. in this case.

OpenStudy (anonymous):

Which in conclusion?

OpenStudy (luigi0210):

means you have to get rid of it by multiplying it out.

OpenStudy (anonymous):

I multiplied and got like 82/90?

OpenStudy (anonymous):

I must be multiplying wrong..or something.

OpenStudy (luigi0210):

Well hold on

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

You type for a long time.-.

OpenStudy (luigi0210):

So I'm gonna take a different approach: 1.) multiply the top and bottom by rad 15: \[\frac{ \sqrt50 }{ 3\sqrt15 }*\frac{ \sqrt15 }{ \sqrt15 }\] A radical 15* radical 15 makes it just a 15, so that leaves us with this: \[\frac{ \sqrt{15*50} }{ 3*(15) }\] 3.) Multiply the radical 15 and 50 giving you 750 and the 3 and 15 giving you 45 \[\frac{ \sqrt{750} }{ 45 }\] 4.)Now break apart the inside if the radical: \[\frac{ \sqrt{25*30} }{ 45 }\] 5.) Since 25 is a perfect square we can take that out as 5. 30 On the other hand just breaks down more into 6*5 so we can leave that inside: \[\frac{ 5 \sqrt{30} }{ 45 }\] 6.) The top and bottom have a common factor of 5 so we can cancel that out leaving us with:

OpenStudy (anonymous):

\[5\sqrt{30}\] ?

OpenStudy (anonymous):

Oh my bad.-.

OpenStudy (anonymous):

\[\sqrt{30} / 9\]

OpenStudy (luigi0210):

Yes :)

OpenStudy (anonymous):

I have two more questions and I just want to be done already .-.

OpenStudy (luigi0210):

xD